Simply and show steps:
square root [144x^10y^12z^18]

Answer by checkley77(12844)

square root [144x^10y^12z^18]
sqrt of 144=12
x^(10/2)=x^5
y^(12/2)=y^6
z^(18/2)=y^9
Thus you have:
12x^5y^6z^9 ans.